What is the area of a rectangle with vertices (2,4),(5,4),(5,10), and (2,10)
Answer by Alan3354(69443)

What is the area of a rectangle with vertices (2,4),(5,4),(5,10), and (2,10)
---------------------
The diff in x is 3 (5-2)
The diff in y is 6 (10-4)
Area = 3*6 = 18 sq units.
----------------------
Another way, if the points are randomly spaced:
A B C D A
2 5 5 +2 +2
4 4 10 10 4 (the +'s are for alignment)
-------------
Add the diagonals from upper left:
2*4 + 5*10 + 5*10 + 2*4 = 116
Add the diagonals from lower left:
4*5 + 4*5 + 10*2 + 10*2 = 80
116 - 80 = 36
Area = 36/2 = 18, same answer
That works for any number of points placed anywhere.